Using the Newton-Raphson method, starting the equation from x0 = 0.7 Solve for 4 iterations?

f(x)=x^{3}+4x^{2}-3  

Using the Newton-Raphson method, starting the equation from x0 = 0.7
Solve for 4 iterations?
Answer #1

Solution: f(x) = x3 + 4x2 - 3 f(x) = 3x2 + 8x Formula used for Newton Raphson method f(x) Xn+1 = xn- f(x) Let the initial r